The school registrar estimates that the dropout rate of freshmen high schools in Mindanao


is 20%. Last year, 45 freshmen from a random sample of 250 Mindanao freshmen high


school withdrew. At š¯›¼ = 0.01, is there enough evidence to reject the registrarā€™s claim?

1
Expert's answer
2022-03-08T04:24:03-0500

"H_0:p=p_0=0.2"

"H_1:p\not=0.2"

Test statistics

"T={\frac {{\frac m n}-p_0} {p_0*(1-p_0)}}*\sqrt{n}" , where m - number of satisfiyung observations, n - sample size

In the given case we have

"T={\frac {{\frac {45} {250}}-0.2} {0.2*0.8}}*\sqrt{250}\approx -1.98"

Since the sample size is big(>30), then it is appropriate to use z-score as critical value C, so, since we run two-tailed test

"P(Z>C)={\frac {0.01} 2}=0.005\implies C=2.58"

Since "T\in(-2.58,2.58)" , then we should accept the null hypothesis and conclude that, based on the data, there is no statisticsal evidence at "\alpha=0.01" to conclude that "p\not=0.2"